# Why do zeroes happen? A model-based approach for demand classification

### v1

@@ 5.5 Results of the experiment
Conventional 0.0035 0.4154 0.6814 0.9535 1.0975 65.2722
Full 0.0030 0.4130 0.6735 0.9247 1.0738 60.0936
Mixture 0.0000 0.2204 0.4642 0.6646 0.8289 125.4443
Category Partial 0.0000 0.2235 0.4677 0.6601 0.8239 68.3509
Category Full 0.0000 0.2221 0.4668 0.6637 0.8251 124.0355
@@
Conventional 0.0002 0.4295 0.6950 0.9620 1.0980 73.5414
Full 0.0004 0.4407 0.7090 0.9596 1.0942 73.6920
Mixture 0.0000 0.3031 0.5770 0.8422 0.9984 79.6121
Category Partial 0.0000 0.3091 0.5826 0.8526 1.0027 76.5965
Category Full 0.0000 0.3090 0.5829 0.8517 1.0019 78.1593
@@
Conventional 0.0000 0.4255 0.6898 0.9625 1.0985 78.3027
Full 0.0000 0.4277 0.6906 0.9552 1.0954 78.3027
Mixture 0.0000 0.3043 0.5762 0.8380 0.9973 75.5905
Category Partial 0.0000 0.3107 0.5867 0.8456 1.0051 75.5905
Table 3: RMSSE values of forecasting approaches with different features on the retail company data. Q1 and Q3 are the first and third quartiles respectively.
@@
    All LightGBM methods are more accurate than the conventional smoothed series (i.e. that ignores stockouts) across all statistics of the RMSSE;
@@
    The split into demand occurrence and demand sizes (“Mixture”) leads to further improvements in terms of RMSSE in comparison with the “Full” model. This improvement is once again observed across the three approaches and we can observe a substantial decreases in the RMSSE;
@@
    The split into Regular/Intermittent categories (“Category Partial”) does not bring substantial value in comparison with the “Mixture” approach:
    • (a)
      In case of the LightGBM, the mean RMSSE decreases, while the median one goes up. The maximum value of the RMSSE decreases, which implies that the approach does not do as big mistakes as the previous one. This is useful in practice where very poor performance of approaches on some observations can raise serious concerns of the data scientist team;
    • (b)
      In case of regression and smoothed series, there is no apparent improvement in terms of mean or quantiles of RMSSE, although the differences do not look substantial;
@@
    The split into finer categories of Regular/Smooth/Lumpy (“Category Full”) does not lead to any consistent noticeable improvements in comparison with the simpler classification (“Category Partial”), although there seems to be a slight reduction in median RMSSE in case of LightGBM and in mean one for the Pooled Regression.
@@
We conducted the Nemenyi test (Demšar, 2006) implemented in the rmcb() function in the greybox package in R (Svetunkov, 2024a) to see whether the differences in the performance of the approaches is statistically significant on the 5% level. Figure 11 depicts the results of this test, showing the average ranks for each of the approaches on the y-axis: the lower the approach is located, the higher rank it has, meaning that it outperforms the others on series-to-series basis more often than the other ones. If the differences in performance between approaches is not significant on the 5% level, the vertical line is drawn, connecting them. If the differences are significant, the dot is placed in the plot. Figure 11 shows performance of approaches per group.
For the LightGBM group, we see that the “Category Full” is outperforming the others in the majority of cases, resulting in the highest rank and being significantly different than the others. “Mixture” and the “Category Partial” are not distinguishable in terms of the statistical significance.
@@
For the Polled Regression, the “Category Partial” is significantly more accurate than the other models. Although in terms of the mean RMSSE, both “Category Partial” and “Category Full” were slightly worse than the “Mixture” model, on the individual level of series, they are doing better.
For the Smoothed series, the “Category Partial” was ranked the highest on average, but it does not differ from the “Mixture” significantly on the 5% level.
@@
Finally, the ordering of “Mixture”, “Full” and “Conventional” is preserved for the three approaches, implying that the respective features indeed bring value.
@@
Summarising the results of this experiment, we see that there is a value in detecting stockouts and including them in a forecasting approach (as long as they are removed from the smoothed series) and that the split into the demand sizes and demand occurrence tends to substantially improve performance of forecasting approaches. The split into the Regular/Intermittent categories tends to further improve performance, but not by the same margin. Finally, the split into finer categories of intermittent demand further reduces significantly forecast errors in case of LightGBM. However, we should note that the dataset at hand was relatively homogenous, and if time series with more different patterns (e.g. a mixture of intermittent and regular with strong seasonality) were present, the split into finer categories could have brought more gains in terms of accuracy.
@@ 6 Conclusions
After that, we developed an Automatic Identification of Demand (AID) approach that automatically detects stockouts and classifies demand into one of the six categories based on AIC of several models underlying each of the types.
@@
Finally, we did an experiment on the real retail data, trying to see whether introducing specific features and using several fundamental modelling principles improves accuracy of several basic forecasting approaches. We found that:
@@
  • 3.
    Applying different approaches to regular and intermittent data leads to further accuracy improvements, although not as substantial;
@@
    The further split into smooth/lumpy leads to slight improvements, but they are not always consistent.
@@
Data availability

### v2

@@ 5.5 Forecasting results
Conventional 0.0026 0.4849 0.7966 1.1242 1.3113 72.1749
Full 0.0056 0.4754 0.7780 1.0907 1.2788 70.4951
Mixture 0.0000 0.4550 0.7574 1.0682 1.2375 81.7721
Category Partial 0.0000 0.4288 0.7125 1.0008 1.1624 62.9427
Category Full 0.0000 0.4296 0.7107 0.9982 1.1600 62.8027
@@
Conventional 0.0007 0.5057 0.8249 1.1669 1.3433 105.3349
Full 0.0004 0.5158 0.8244 1.1492 1.2852 251.5448
Mixture 0.0000 0.4597 0.7657 1.0766 1.2486 82.0593
Category Partial 0.0000 0.4255 0.7125 1.0016 1.1623 78.1823
Category Full 0.0000 0.4275 0.7142 1.0006 1.1622 79.5183
@@
Conventional 0.0000 0.5023 0.8148 1.1481 1.3198 82.2092
Full 0.0000 0.5042 0.8148 1.1427 1.3091 82.2092
Mixture 0.0000 0.4565 0.7600 1.0686 1.2382 82.0491
Table 3: RMSSE values of forecasting approaches with different features on the retail company data. Q1 and Q3 are the first and third quartiles respectively. The lowest error measures in each category are shown in boldface, the lowest overall are in boldface and italic.
@@
    All LightGBM methods are more accurate than the conventional smoothed series (those that ignore stockouts) across all statistics of the RMSSE;
@@
    The split into demand occurrence and demand sizes (“Mixture”) leads to further improvements in terms of RMSSE in comparison with the “Full” model for all approaches with a substantial decreases in RMSSE;
@@
    The introduction of the Category feature (which only has “regular/intermittent” ones) further improves the accuracy both in case of the LightGBM and regression;
@@
    The split into regular/smooth intermittent/lumpy intermittent leads to further improvement in accuracy (“Category Full”), except for the median RMSSE in case of regression.
@@
We conducted the Nemenyi test (Demsar2006) implemented in the rmcb() function in the greybox package in R (Svetunkovgreybox) to see whether the differences in the performance of the approaches is statistically significant on the 5% level. Figure 12 depicts the results of this test, showing the average ranks for each of the approaches on the y-axis: the lower the approach is located, the higher rank it has, meaning that it outperforms the others on series-to-series basis more often than the other ones. If the differences in performance between approaches is not significant on the 5% level, the vertical line is drawn, connecting them. If the differences are significant, the dot is placed in the plot. Figure 12 shows performance of approaches per group.
For the LightGBM group, we see that the “Category Full” outperforms the others in the majority of cases, resulting in the highest rank. In fact, we see that the ranking observed in Table 3 is preserved, and the main observations discussed above stay.
@@
For the Pooled Regression, the “Category Partial” dominates the others in terms of accuracy. Although in terms of the mean RMSSE, “Category Full” was slightly better than the “Category Partial”, on the individual level of series, it does better.
For the Smoothed series, the “Mixture” was ranked the highest on average, outperforming the others significantly on the 5% level. Note that the “Conventional” is not statistically different from the “Full” category in this case. This shows that the main benefit comes not from the smoothed series, but rather from the stockout feature itself, which cannot be included in case of just Smoothed time series.
@@
Most importantly, the ordering of “Category”, “Mixture”, “Full”, “Conventional” is roughly preserved for the three approaches, implying that the respective features indeed bring value.
@@
Summarising the results of this experiment, we see that there is a value in detecting stockouts and including them in a forecasting approach (as long as they are removed from the smoothed series) and that the split into the demand sizes and demand occurrence tends to substantially improve performance of forecasting approaches. The split into the Regular/Intermittent categories, which is done automatically after detecting stockouts, tends to further improve performance. Finally, the split into finer categories of demand further reduces significantly forecast errors in case of LightGBM and helps substantially in the pooled regression. The latter can be attributed to the fact that regression is a linear model, and a split into finer categories makes it more flexible.
5.6 Inventory simulation
To better understand whether the usage of the AID approach translates to better inventory decisions, we setup a simple inventory simulation for the available data. We did not have the information about the prices of products, so we simplified the whole setup to the following:
  • 1.
    we used the order-up-to-level policy,
  • 2.
    the lead time was set to one week, because for the majority of products it was on average close to one,
  • 3.
    we assume that all the products expire after one week. While this is a strong assumption for so many SKUs, we did not have the information about the expiration dates, and we argue that this assumption does not violate the main findings of the experiment,
  • 4.
    we do the simulation with rolling origin for two steps, not updating the features (i.e. not recalculating the smoothed series),
  • 5.
    we used service levels of 90%, 95% and 99% recording the achieved service level, scaled lost sales and inventory on hand similarly to how it was done by Kourentzes2019.
To get the safety stock levels, we used a rather simple approach, relying on the empirical quantile calculation, similar to how it was done by Lee2014 and then adapted to inventory management by Trapero2019 (nowadays approaches like this are sometimes called “conformal prediction”). In nutshell, the approach can be summarised in following steps:
  1. 1.
    Take a LightGBM approach used in the previous section;
  2. 2.
    Apply it to the training data to get the fitted values $\hat{y}_{t}$;
  3. 3.
    Generate forecast errors $e_{t}=y_{t}-\hat{y}_{t}$;
  4. 4.
    Scale forecast errors by the standard deviation of each separate time series;
  5. 5.
    Extract the necessary quantiles of the forecast errors;
  6. 6.
    Scale back produced quantiles for each series by multiplying them by standard deviations;
  7. 7.
    Produce point forecasts for each time series (this was done in Subsection 5.5);
  8. 8.
    Add quantiles to each of the point forecasts to get the safety stock levels;
  9. 9.
    Round up the final values, because we deal with integer-valued data.
The advantage of this simple approach is that it is fast and does not require retraining any of the approaches. The disadvantage is that it relies on the in-sample forecast errors, which might lead to potential overfitting and thus lead to lower quantiles than needed. But we argue that it should suffice for the demonstration purposes: if we see the same sequence in improvement for the approaches as we saw in Subsection 5.5, this would mean that AID works for this dataset. If some of the aspects in this inventory simulation are improved, we argue that the sequence should stay the same.
In case of the mixture approaches (this applies to “Mixture”, “Category Partial” and “Category Full”), the target service level was amended for each individual observation depending on the predicted probability of occurrence. This was done similar to how Svetunkov2023b did it in Subsection 4.3 for the iETS model, using the formula:
$F_{z}(z_{t+h}\leq Q)=\frac{F_{y}(y_{t+h}\leq Q)-(1-\hat{p}_{t+h|t})}{\hat{p}_{t+h|t}},$ (3)
where $F_{y}(y_{t+h}\leq Q)$ is the target service level for the actual value $y_{t+h}$, $\hat{p}_{t+h|t}$ is the probability of occurrence conditional on the information on the observation $t$, $F_{z}(z_{t+h}\leq Q)$ is the confidence level calculated for the demand sizes $z_{t+h}$, and $Q$ is the quantile of interest.
We decided to focus on LightGBM because it produce the most accurate point forecasts in the previous subsection. However, we checked the inventory implications for other approaches and found similar results in terms of their ranking.
5.7 Inventory simulation results
Figure 13: Lost sales vs inventory on hand and target service level vs the service level deviation. Origin 1.
Lost sales and inventory on hand were scaled by dividing them by the in-sample standard deviation of the data. We also calculated “Service level deviation” by subtracting the achieved service level from the target one. Figure 13 shows these measures in one plot for the first step. For the lost sales vs inventory on hand image, the best performing approach is the one that is closest to the origin. For the service level image, the perfect approach should lie on the horizontal line going through zero (implying that the achieved service level equals the target one).
As we can see from the first image in Figure 13, the “Conventional” approach leads to highest lost sales and inventory on hand in comparison with the other approaches. The “Full” approach which relies on stockouts detection improves upon it, leading to slightly higher inventory on hand but with the reduced lost sales. The categorisation approaches (which split the data into regular/intermittent and to regular/smooth/lumpy), together with the “Mixture” one bring further improvements, mainly reducing the lost sales. The three perform very similar, and their performance is indistinguishable. All three approaches produce the curve that is the closest to the origin, although it leads to slightly higher lost sales (at the cost of lower inventory on hand) for the low service level of 90% than in case of the “Full” approach. When the service level is the highest (99%), it seems to produce the best combination of the lost sales and inventory on hand.
Figure 14: Lost sales vs inventory on hand and target service level vs the service level deviation. Origin 2.
When it comes to the achieved vs target service level (the second image in Figure 13), we see that the conventional approach consistently underperforms, containing fewer than expected observations. Moving to the “Full” approach, the difference becomes apparent: the improvement is substantial and the SL deviation becomes close to zero (being slightly below, implying that the the achieved SL is less than 1% lower than the target). This once again demonstrates the importance of treating the stockouts correctly. Finally, moving to the “Mixture”, “Category Partial” and “Category Full” approaches, the lines lie consistently slightly above zero, meaning that the achieved SL is higher than the target (the deviation is less than 1%).
The whole situation changes for the second week as shown in Figure 14, although the ranking stays the same, which shows that the AID principles work in terms of inventory management for the dataset. The three approaches (“Mixture”, “Category Partial” and “Category Full”) still perform similar, leading to lower lost sales although at the cost of inventory on hand (in comparison with the other approaches). Furthermore, the three approaches lead to lower SL deviation than the other two approaches, although the difference between them and the “Full” one seems to be negligible.
As a summary from this experiment, we found that the inclusion of stockouts feature improves substantially the performance of LightGBM in terms of inventory. Furthermore, the split into demand occurrence and demand sizes leads to further improvements. However, the AID classification itself does not bring any benefits in terms of inventory performance in comparison with the “Mixture” approach.
Finally, we also included the count/fractional feature and experimented with different ways for generating quantiles based on this split (using different distributions and quantile generation approaches), but we did not find any improvement in terms of inventory performance in comparison with the “Mixture” approach. So, we argue that the split into count/fractional, while being theoretically appealing, is not practically useful.
@@ 6 Conclusions
After that, we developed an Automatic Identification of Demand (AID) approach that automatically detects stockouts and first classifies demand into regular or intermittent (depending on the presence of naturally occuring zeroes), and then does a fine split to one of the six categories based on AIC of several models underlying each of the types.
@@
Furthermore, we did an experiment on the real retail data, trying to see whether introducing specific features and using several fundamental modelling principles improves accuracy of several basic forecasting approaches. We found that:
@@
    The stockouts detection should be done for both the training and the test sets. If the series with stockouts are not removed from the test set, the forecasts would be evaluated incorrectly (i.e. showing which of the methods does better at predicting sales instead of demand);
  • 3.
@@
    Using the feature for regular/intermittent demand improves the forecasting accuracy, but does not seem to impact the inventory performance. Note that this separation is straightforward in AID: if after removing the stockouts, there are some zeroes left, the demand is identified as intermittent;
  • 5.
    The further split into smooth/lumpy leads to slight improvements in terms of accuracy, without a substantial impact on the inventory;
  • 6.
    The split into count/fractional demand does not bring value in terms of forecasting accuracy or inventory performance.
@@
Finally, we conducted additional experiment on the M5 daily data (Makridakis2022), setting the forecasting horizon to 14 days, applying the same approaches as in the case study to test the robustness of the AID approach. We found that the main principles discussed in this paper hold there as well.
In terms of the limitations, we note that other information criteria or cross validated error measures could be used instead of the AIC in the AID approach. We used it for simplicity, but an investigation of the impact of information criteria on the overall performance of the algorithm could be done in a future work.
We should also remark that we did not have the records of stockouts from the company to test whether the detection mechanism works well and in how many cases it makes a mistake. It would be interesting to explore this element in future work. But the more important message from the paper is that whatever detection mechanism you use for stockouts, they need to be treated and included in your model, because this becomes essential in terms of forecasting accuracy and inventory management.
Data availability and competing interests